use source transformation to find the current io.

2.3 14л бо 2 mA (T (1) 2.3ik 2014 Л. 0. GMA
Answer #1

Given circuit; A 2.3ka M To 0.6mA 2 MA {2.7k Zike & Source Transformation. By using R ML R + V - IR * I R=2.7k2 M -3 3 2.7k2Circuit can be redrawn as 2.7 km 2.3ka 1 kr 0- A Io + + 0.6V 5.4v Р 5.40 0.6 + Io +2.3 Io + 2.7 Io + I. = - 0.6 - 5 - 4 (1 +